# August 2025 IPMU25-0038

### v1

@@ 5 More on SR approximation in Jordan frame
$\alpha_{s}\approx-2\epsilon_{2}\epsilon_{3}\approx-2\zeta_{2}\zeta_{3}$ (83)
@@
$\alpha_{t}\approx-12\epsilon_{1}^{2}\epsilon_{2}\approx-3\zeta_{1}^{2}\zeta_{2}~,$ (84)
@@
$A_{s}\approx{{H^{2}}\over{\pi^{2}rF_{,R}}}\approx{{{\left(2F_{,R}R-F\right)^{4}}\over{2592\,{\pi}^{2}F_{,R}^{2}\left(F_{,R}R-F\right)\left(F_{,R}R-2F\right)^{2}}}}\,.$ (85)
In terms of the functions $U(\sigma)$ and $V(\sigma)$, Eqs. (70) and (72) take the form
${R\,}^{\prime}(\sigma)\approx{{2M_{\mathrm{Pl}}^{4}UV_{\mathrm{eff},\sigma}}\over{M_{\mathrm{Pl}}^{4}U_{,\sigma}V_{\mathrm{eff},\sigma}+4AU^{2}V_{\mathrm{eff}}}}={}-{{2U\left(2VU_{,\sigma}-V_{,\sigma}U\right)}\over{U_{,\sigma}\left(VU_{,\sigma}+V_{,\sigma}U\right)}}$ (86)
@@
$H^{2}(\sigma)\approx{{\left(VU_{,\sigma}+V_{,\sigma}U\right)^{2}}\over{54VUU_{,\sigma}^{2}}}\,$~, (87)
@@ 6 Starobinsky model in Jordan frame
It is instructive to apply the equations in the preceding Sections to the well-known case of the Starobinsky inflation model described by Eq. (1), without resorting to the Einstein frame or the corresponding inflaton potential (2).
@@
$F(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6\,m^{2}}}\right)~,$ (88)
@@
$H^{2}\approx{{\left(2{m}^{2}+R\right)^{2}}\over{12\left(3\,{m}^{2}+R\right)}}\approx{{R}\over{12}}$ (89)
@@
${R\,}^{\prime}\approx{{4{m}^{2}\left(3{m}^{2}+R\right)}\over{2m^{2}+R}}\approx 4m^{2}~,$ (90)
@@
After solving Eq. (90), we get the running e-folds as
$N={{R}\over{4m^{2}}}-{{1}\over{4}}\ln{{3m^{2}+R}\over{R_{0}}}\,$, (91)
@@
$\zeta_{1}(z)={}-4{{{z}\over{2z+1}}}\approx{}-4z~,$ (93)
$\zeta_{2}(z)=4{{{z\left(3z+1\right)}\over{\left(2z+1\right)^{2}}}}\approx 4z~,$ (94)
$\zeta_{3}(z)=4{{{z\left(4\,z+1\right)}\over{\left(2\,z+1\right)^{2}}}}\approx 4z~,$ (95)
$\varepsilon_{1}(z)=2\,{{{z\left(4z+1\right)}\over{\left(2z+1\right)^{2}}}}\approx 2z~,$ (96)
$\varepsilon_{2}(z)=4{{{z\left(6z+1\right)\left(3z+1\right)}\over{\left(2z+1\right)^{2}\left(4z+1\right)}}}\approx 4z~,$ (97)
@@
$\varepsilon_{3}(z)=4{{{z\left(108{z}^{3}+74\,{z}^{2}+16z+1\right)}\over{\left(2z+1\right)^{2}\left(4z+1\right)\left(6z+1\right)}}}\approx 4z\,.$ (98)
Hence, in the leading order, we get from Eqs. (93)–(97) that
$\zeta_{1}\approx-2\varepsilon_{1}~,\quad\zeta_{2}\approx\varepsilon_{2}\approx 2\varepsilon_{1},\quad\zeta_{3}\approx\varepsilon_{3}\approx 2\varepsilon_{1}~.$ (99)
Using Eqs. (81) and (82), we also obtain
$n_{s}\approx 1-2\zeta_{2}=1-8z{{\left(3z+1\right)}\over{\left(2z+1\right)^{2}}}$ (100)
@@
$r\approx 12\zeta_{1}^{2}={{192z^{2}}\over{\left(2z+1\right)^{2}}}\,.$ (101)
Hence, the leading terms in Eqs. (54)–(57) are given by
$r\approx 48\varepsilon_{1}^{2}+{\cal O}(\varepsilon_{1}^{3})~,\quad n_{s}\approx 1-4\varepsilon_{1}+{\cal O}(\varepsilon_{1}^{2})~,$ (102)
$\alpha_{s}\approx{}-8\varepsilon_{1}^{2}+{\cal O}(\varepsilon_{1}^{3})~,\quad\alpha_{t}\approx{}-24\varepsilon_{1}^{3}+{\cal O}(\varepsilon_{1}^{4})~.$ (103)
@@
$R_{\rm end}\approx(\sqrt{5}-1)m^{2}\approx 1.2m^{2}~.$ (104)
@@
Table 1: The values of $R_{\rm in}$, $r$, and $N_{*}$ in the Starobinsky model.
@@
In the leading order of the SR approximation, we find from Eq. (91) that
$N_{*}\approx{{1}\over{4z}}\approx{{1}\over{2\varepsilon_{1}}}~.$ (105)
Then Eq. (102) reproduces the well-known results for the Starobinsky model,
$n_{s}=1-{{2}\over{N_{*}}}+{\cal{O}}(N_{*}^{-2}),\qquad r={{12}\over{N_{*}^{2}}}+{\cal{O}}(N_{*}^{-3})~,$ (106)
@@
$\alpha_{s}={}-{{2}\over{N^{2}_{*}}}+{\cal{O}}(N_{*}^{-3}),\qquad\alpha_{t}={}-{{3}\over{N_{*}^{3}}}+{\cal{O}}(N_{*}^{-4})~.$ (107)
@@ 7 ACT versus Planck
$n_{s}=0.9651\pm 0.0044\,,\quad\alpha_{s}\equiv\mathrm{d}\,n_{s}/\mathrm{d}\ln k=-0.0069\pm 0.0069~,$ (108)
@@
$n_{s}=0.9743\pm 0.0034\,,\quad\alpha_{s}\equiv\mathrm{d}\,n_{s}/\mathrm{d}\ln k=0.0062\pm 0.0052~.$ (109)
@@
$A_{s}=(2.10\pm 0.03)\times 10^{-9}\qquad{\rm and}\qquad r<0.028~.$ (110)
@@
Nevertheless, it makes sense to confront the existing single-field models of inflation with the ACT constraints given above and ask for possible cure. It is not difficult to increase the value of $n_{s}$, while keeping the desired duration of inflation within $50\div 60$ e-folds, by adding extra terms to the known models and adjusting extra parameters to get an ACT-conform inflation potential during inflation in the SR approximation, see e.g., Refs. [18, 19, 41, 42, 43, 44, 45, 46, 47] in the Einstein frame. However, it is more difficult to change the sign of $\alpha_{s}$ in the SR approximation. Actually, the majority of single-field inflation models predict negative values of $\alpha_{s}$ [1] and, hence, upward bending of the inflaton potential. The ACT-favored downward bending of the inflaton potential for the higher values of the inflaton field before inflation is not an issue to be worried about because those values are beyond the scope of inflation.
As regards the Starobinsky inflation model in the Jordan frame, an obvious phenomenological approach to meet the ACT constraints is adding the higher order terms with respect to the spacetime scalar curvature in the context of $F(R)$ gravity. Alternatively, from the fundamental viewpoint, string theory corrections to the gravitational effective action may do the same job [38]. As was demonstrated in Ref. [18], it is possible to increase the value of $n_{s}$ in Starobinsky inflation to the ACT-favored region, while having small negative values of $\alpha_{s}$, by adding the $R^{3}$- or $R^{4}$- corrections to the action (1).
@@
8 New modified model of Starobinsky inflation
Since the ACT observations have merely $\sim 2\sigma$ tension with the predictions of the Starobinsky model, it is reasonable to study small deformations of the model in the context $F(R)$ gravity for the limited range of the spacetime scalar curvature $R$ relevant to inflation. Since those modifications are small and conservative, they belong to the same universality class of inflationary models as the Starobinsky model itself, so that all such models can be referred to as Starobinsky inflation where the $R^{2}$-term is responsible for flatness of the inflaton potential.
@@ 8 New modified model of Starobinsky inflation
$F_{3}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6m^{2}}}+{{\delta_{3}R^{3}}\over{36m^{4}}}\right)$ (111)
@@
$F_{4}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6m^{2}}}+{{\delta_{4}R^{4}}\over{48m^{6}}}\right)$ (112)
@@
$F_{5}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{1}\over{6m^{2}}}R^{2}+{{c_{3}}\over{m^{4}}}R^{3}+{{c_{4}}\over{m^{6}}}R^{4}+{{c_{5}}\over{m^{8}}}R^{5}\right)~,$ (113)
@@
$\zeta_{1}={}-{{4\left({z}^{4}-c_{3}{z}^{2}-2c_{4}z-3c_{5}\right)}\over{2{z}^{4}+{z}^{3}+10c_{3}{z}^{2}+14c_{4}z+18c_{5}}}~.$ (114)
The observational bound $r<0.028$ and Eq. (81) also imply
$|\zeta_{1_{in}}|<0.048~.$ (115)
Hence, any term except the $R^{2}$ in Eq. (113) should not dominate, all the parameters $c_{i}$ should be small, $\left|c_{i}\right|\ll 1$, whereas a small positive value of $c_{5}$ can be responsible for changing the sign of $\zeta_{1}$.
Using the SR approximation formulae (81), (82) and (83), the values of $c_{i}$ can be chosen to meet the ACT favored values of $n_{s}$ and $\alpha_{s}$. Our results are summarized in Table 2. Those results are robust against small changes in the values of $c_{i}$, while $R_{in}$ can be adjusted to keep $N_{*}$ in the desired range.
The stability conditions (6) are satisfied for all positive $R$. The effective potential $V_{eff}$ has a maximum at $R>150\,m^{2}$, so that $R(t)$ is a monotonically decreasing function during inflation. The values of $n_{s}$, $r$, and $\alpha_{s}$ are independent upon the value of $m$ that is fixed by the observed value of $A_{s}$.

### v2

@@ 5 More on SR approximation in Jordan frame
$\alpha_{s}\approx-2\epsilon_{2}\epsilon_{3}\approx-2\zeta_{2}\zeta_{3}$ (5.17)
@@
$\alpha_{t}\approx-12\epsilon_{1}^{2}\epsilon_{2}\approx-3\zeta_{1}^{2}\zeta_{2}~,$ (5.18)
@@
$A_{s}\approx{{H^{2}}\over{\pi^{2}rF_{,R}}}\approx{{{\left(2F_{,R}R-F\right)^{4}}\over{2592\,{\pi}^{2}F_{,R}^{2}\left(F_{,R}R-F\right)\left(F_{,R}R-2F\right)^{2}}}}\,.$ (5.19)
In terms of the functions $U(\sigma)$ and $V(\sigma)$, Eqs. (5.4) and (5.6) take the form
${R\,}^{\prime}(\sigma)\approx{{2M_{\mathrm{Pl}}^{4}UV_{\mathrm{eff},\sigma}}\over{M_{\mathrm{Pl}}^{4}U_{,\sigma}V_{\mathrm{eff},\sigma}+4AU^{2}V_{\mathrm{eff}}}}={}-{{2U\left(2VU_{,\sigma}-V_{,\sigma}U\right)}\over{U_{,\sigma}\left(VU_{,\sigma}+V_{,\sigma}U\right)}}$ (5.20)
@@
$H^{2}(\sigma)\approx{{\left(VU_{,\sigma}+V_{,\sigma}U\right)^{2}}\over{54VUU_{,\sigma}^{2}}}\,$~, (5.21)
@@ 6 Starobinsky model in Jordan frame
It is instructive to apply the equations, obtained in the preceding sections, to the well-known case of the Starobinsky inflation model described by Eq. (1.1), without resorting to the Einstein frame or the corresponding inflaton potential (1.2).
@@
$F(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6\,m^{2}}}\right)~,$ (6.1)
@@
$H^{2}\approx{{\left(2{m}^{2}+R\right)^{2}}\over{12\left(3\,{m}^{2}+R\right)}}\approx{{R}\over{12}}$ (6.2)
@@
${R\,}^{\prime}\approx{{4{m}^{2}\left(3{m}^{2}+R\right)}\over{2m^{2}+R}}\approx 4m^{2}~,$ (6.3)
@@
After solving Eq. (6.3), we get the running e-folds as
$N={{R}\over{4m^{2}}}-{{1}\over{4}}\ln{{3m^{2}+R}\over{R_{0}}}\,$, (6.4)
@@
$\zeta_{1}(z)={}-4{{{z}\over{2z+1}}}\approx{}-4z~,$ (6.6)
$\zeta_{2}(z)=4{{{z\left(3z+1\right)}\over{\left(2z+1\right)^{2}}}}\approx 4z~,$ (6.7)
$\zeta_{3}(z)=4{{{z\left(4\,z+1\right)}\over{\left(2\,z+1\right)^{2}}}}\approx 4z~,$ (6.8)
$\varepsilon_{1}(z)=2\,{{{z\left(4z+1\right)}\over{\left(2z+1\right)^{2}}}}\approx 2z~,$ (6.9)
$\varepsilon_{2}(z)=4{{{z\left(6z+1\right)\left(3z+1\right)}\over{\left(2z+1\right)^{2}\left(4z+1\right)}}}\approx 4z~,$ (6.10)
@@
$\varepsilon_{3}(z)=4{{{z\left(108{z}^{3}+74\,{z}^{2}+16z+1\right)}\over{\left(2z+1\right)^{2}\left(4z+1\right)\left(6z+1\right)}}}\approx 4z\,.$ (6.11)
Hence, in the leading order, we get from Eqs. (6.6)–(6.10) that
$\zeta_{1}\approx-2\varepsilon_{1}~,\quad\zeta_{2}\approx\varepsilon_{2}\approx 2\varepsilon_{1},\quad\zeta_{3}\approx\varepsilon_{3}\approx 2\varepsilon_{1}~.$ (6.12)
Using Eqs. (5.15) and (5.16), we also obtain
$n_{s}\approx 1-2\zeta_{2}=1-8z{{\left(3z+1\right)}\over{\left(2z+1\right)^{2}}}$ (6.13)
@@
$r\approx 12\zeta_{1}^{2}={{192z^{2}}\over{\left(2z+1\right)^{2}}}\,.$ (6.14)
Hence, the leading terms in Eqs. (4.21)–(4.24) are given by
$r\approx 48\varepsilon_{1}^{2}+{\cal O}(\varepsilon_{1}^{3})~,\quad n_{s}\approx 1-4\varepsilon_{1}+{\cal O}(\varepsilon_{1}^{2})~,$ (6.15)
$\alpha_{s}\approx{}-8\varepsilon_{1}^{2}+{\cal O}(\varepsilon_{1}^{3})~,\quad\alpha_{t}\approx{}-24\varepsilon_{1}^{3}+{\cal O}(\varepsilon_{1}^{4})~.$ (6.16)
@@
$R_{\rm end}\approx(\sqrt{5}-1)m^{2}\approx 1.2m^{2}~.$ (6.17)
@@
Table 1: The values of $n_{s}$, $r$, $R_{\rm in}$, and $N_{*}$ in the Starobinsky model.
@@
In the leading order of the SR approximation, we find from Eq. (6.4) that
$N_{*}\approx{{1}\over{4z}}\approx{{1}\over{2\varepsilon_{1}}}~.$ (6.18)
Then Eq. (6.15) reproduces the well-known results for the Starobinsky model,
$n_{s}=1-{{2}\over{N_{*}}}+{\cal{O}}\left(N_{*}^{-2}\right)\,,\qquad r={{12}\over{N_{*}^{2}}}+{\cal{O}}\left(N_{*}^{-3}\right)\,$, (6.19)
@@
$\alpha_{s}={}-{{2}\over{N^{2}_{*}}}+{\cal{O}}\left(N_{*}^{-3}\right)\,,\qquad\alpha_{t}={}-{{3}\over{N_{*}^{3}}}+{\cal{O}}\left(N_{*}^{-4}\right)\,.$ (6.20)
@@ 7 ACT versus Planck
$n_{s}=0.9651\pm 0.0044\,,\quad\alpha_{s}=-0.0069\pm 0.0069~,$ (7.1)
@@
$n_{s}=0.9743\pm 0.0034\,,\quad\alpha_{s}=0.0062\pm 0.0052~.$ (7.2)
@@
$A_{s}=(2.10\pm 0.03)\times 10^{-9}\qquad{\rm and}\qquad r<0.028~.$ (7.3)
@@
Nevertheless, it makes sense to confront the existing single-field models of inflation with the ACT constraints given above and ask for a possible cure. It is not difficult to increase the value of $n_{s}$, while keeping the desired duration of inflation within $50\div 60$ e-folds, by adding extra terms to the known models and adjusting extra parameters to obtain an ACT-conforming inflaton potential during inflation in the SR approximation in the Einstein frame, see e.g., Refs. [18, 19, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. However, it is more difficult to change the sign of $\alpha_{s}$ in the SR approximation. Actually, the majority of single-field inflation models predict negative values of $\alpha_{s}$ [3] and, hence, upward bending of the inflaton potential. The ACT-favored downward bending of the inflaton potential for the higher values of the inflaton field before inflation is not an issue to be worried about because those values are beyond the scope of inflation.
As regards the Starobinsky inflation model in the Jordan frame, an obvious phenomenological approach to meet the ACT constraints is adding the higher order terms with respect to the spacetime scalar curvature in the context of $F(R)$ gravity. Alternatively, from the fundamental viewpoint, string theory corrections to the gravitational effective action may do the same job [41]. As was demonstrated in Ref. [18], it is possible to increase the value of $n_{s}$ in Starobinsky inflation to the ACT-favored region, while having small negative values of $\alpha_{s}$, by adding the $R^{3}$- or $R^{4}$- corrections to the action (1.1).
@@
8 ACT/DESI-improved model of Starobinsky inflation
Since the ACT observations have merely $\sim 2\sigma$ tension with the predictions of the Starobinsky model, it is reasonable to study small deformations of the model in the context of $F(R)$ gravity for the limited range of the spacetime scalar curvature $R$ relevant to inflation. Since those modifications are small and conservative, they belong to the same universality class of inflationary models as the Starobinsky model itself, so that all such models can be referred to as Starobinsky inflation where the $R^{2}$-term is responsible for flatness of the inflaton potential during inflation.
@@ 8 ACT/DESI-improved model of Starobinsky inflation
$F_{3}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6m^{2}}}+{{\delta_{3}R^{3}}\over{36m^{4}}}\right)$ (8.1)
@@
$F_{4}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{R^{2}}\over{6m^{2}}}+{{\delta_{4}R^{4}}\over{48m^{6}}}\right)$ (8.2)
@@
$F_{5}(R)={{M^{2}_{\rm Pl}}\over{2}}\left(R+{{1}\over{6m^{2}}}R^{2}+{{c_{3}}\over{m^{4}}}R^{3}+{{c_{4}}\over{m^{6}}}R^{4}+{{c_{5}}\over{m^{8}}}R^{5}\right)~,$ (8.3)
@@
$\zeta_{1}={}-{{4\left({z}^{4}-c_{3}{z}^{2}-2c_{4}z-3c_{5}\right)}\over{2{z}^{4}+{z}^{3}+10c_{3}{z}^{2}+14c_{4}z+18c_{5}}}~.$ (8.4)
The observational bound $r<0.028$ and Eq. (5.15) also imply
$|\zeta_{1_{in}}|<0.048~.$ (8.5)
Hence, any term except the $R^{2}$ in Eq. (8.3) should not dominate, all the parameters $c_{i}$ should be small, $\left|c_{i}\right|\ll 1$, whereas a small positive value of $c_{5}$ can be responsible for changing the sign of $\zeta_{1}$.
Using the SR approximation formulae (5.15), (5.16), and (5.17), the values of $c_{i}$ can be chosen to meet the ACT favored values of $n_{s}$ and $\alpha_{s}$. Our results are summarized in Table 2. Those results are robust against small changes in the values of $c_{i}$, while $R_{in}$ can be adjusted to keep $N_{*}$ within the desired range.
The stability conditions (2.4) are satisfied for all positive $R$. The effective potential $V_{eff}$ has a maximum at $R>150\,m^{2}$, so that $R(t)$ is a monotonically decreasing function during inflation. The values of $n_{s}$, $r$, and $\alpha_{s}$ are independent upon the value of $m$ that is fixed by the observed value of $A_{s}$.
@@
9 Generalisations of the model for smaller scales
The curvature corrections to the Starobinsky model (1.1), given by Eq. (8.3) with the coefficients from Table 2, do not affect physics at other scales smaller than the scale of inflation, where those corrections are negligible. In particular, the transition to reheating and the radiation-dominated era in our model (8.3) are the same as those in the original Starobinsky model (1.1).
Cosmological inflation in the early Universe has many similarities with dark energy (DE) in the current Universe because both describe the Universe expansion with acceleration. It is possible to (phenomenologically) unify both descriptions in $F(R)$ gravity alone, though finding a viable $F(R)$-function obeying all physical and observational constraints is non-trivial. A solution to this problem was proposed by Appleby, Battye and Starobinsky (ABS) in Ref. [56], without destroying inflation and the standard Big Bang cosmology. The $F(R)$-gravity function proposed in Ref. [56] reads
[…]
where $m$ is the Starobinsky mass, the ABS parameter $E_{AB}$ is defined by
$E_{AB}={{R_{0}}\over{2g\ln(1+e^{2b})}}~,$ (9.2)
while the other parameters $g$ and $b$ of the order one define the shape of the corresponding scalar potential. The new physical scale in the ABS model is given by the parameter $R_{0}$, whose value in the case of the present DE is determined by the current Hubble parameter, $R_{0}\sim H^{2}_{0}$.
Just adding our corrections beyond the $R^{2}$-term from Eq. (8.3) to the ABS-proposed $F(R)$-function (9.1) defines the unified $F(R)$-gravity model of inflation (or primordial DE) and the current DE, this time in agreement with ACT also, leading to a single scalar field in the Einstein frame. The corresponding scalar potential is only locally defined and cannot be computed analytically, though numerical calculations are possible. In contrast, in our approach in the Jordan frame, all our $F(R)$-gravity functions in use are elementary, well-defined and smooth, while there is no necessity to go to the Einstein frame in order to get physical predictions. The $F(R)$-function of the unified model reads
[…]
${{c_{3}}\over{m^{4}}}R^{3}+{{c_{4}}\over{m^{6}}}R^{4}+{{c_{5}}\over{m^{8}}}R^{5}~.$ (9.3)
Moreover, the same $F$-function (9.3) but with the high scale $R_{0}\sim m^{2}$ under the scale of inflation can be used to describe efficient PBHs production after Starobinsky inflation, as was demonstrated in Refs. [34, 35] after fine-tuning the parameters as $g\approx 2.25$ and $b\approx 2.89$. The corresponding scalar potential has two plateaus with a nearly-inflection point between them, which implies the existence of an USR regime. The produced PBHs have masses in the asteroid-size range between $10^{16}$ g and $10^{20}$ g, beyond the Hawking evaporation limit, so that they can be considered as viable candidates for the current dark matter. The PBHs-induced gravitational waves have the frequencies about $10^{-2}$ Hz, which makes them detectable by the future space-based gravitational interferometers such as LISA [57] and DECIGO [58], while the quantum (loop) corrections to the power spectrum are small, see Refs. [34, 35] for a derivation of those results.
It is also possible to add the ABS terms to our model (8.3) twice with proper values of the parameters but the different scales corresponding to the PBHs scale and the DE scale, respectively, in order to simultaneously describe inflation, PBHs production and current DE. This shows the power of our minimal approach based on $F(R)$ gravity in the Jordan frame without adding extra matter. Coupling matter scalars to $F(R)$ gravity for describing ACT-consistent inflation with PBHs production was studied in Refs. [50, 59].
Finally, as was noticed in Ref. [56], the number of e-folds $N$ for inflation often increases in the unified $F(R)$ gravity models, while in the case of Starobinsky inflation $N$ may increase up to $70$. In turn, it leads to an ACT-favoured increase of the scalar tilt $n_{s}$. Though it is not enough to get the desired ACT range of $n_{s}$, it implies less fine-tuning of the coefficients $c_{i}$ in Eq. (9.3) that makes our model more robust.